# Bias Injection Attacks on RAG Databases and Sanitization Defenses

### v1

@@ 4.4 Impact on the LLM Answers
Unattacked average absolute PS (Avg. ||PS||): This value, as reported in Tables IV and V, measures the original bias level in unattacked cases. It also indicates whether the defender misses benign passages with strong perspectives, resulting in more neutral and ambiguous content. BRRA, MMR, and SMART yield lower Avg. ||PS|| than No Def. for both retrieved passages and answers. This is because the diversity mechanism may attenuate strong perspectives.
PS shift caused by the attack: The attacker intends to make the PS of the retrieved passages biased toward its desired direction (i.e., ascending direction for positive attack and descending direction for negative attack). We report the average PS shift (see Eq. (4)) of the top-5 retrieved passages. The bias injection attack causes the retrievers to retrieve biased passages, and their average PS shift increases as more adversarial passages are injected (see Table IV). This demonstrates the vulnerability of retrieval systems to deliberate bias manipulation through adversarial passages.
We also report on the average PS shift of the generated answers to every question in the corpus. As listed in Table V, the bias injection attack successfully induces PS shifts in the answers, and the magnitude of the shift increases with the number of injected adversarial passages. Moreover, the generated answers are positively correlated with the retrieved passages in terms of PS values. Fig. 4 quantifies such correlations.
Retrieval bias propagates through the generation process. LLM models fail to mitigate the injected biases during answer synthesis, even when equipped with critical thinking instructions in the prompt (see Prompt-Generation in Appendix B for the prompt used).
Vulnerability across models and datasets: A natural question arises: Does the impact of bias injection attacks remain if we switch to a different—or larger—LLM models? To answer this, we evaluate multiple LLMs using identical retrieval contexts, thereby isolating the model’s role in absorbing contextual bias. Table VI presents representative results for the undefended retriever on WIKI-BALANCE.
Our results show that all three LLMs exhibit comparable PS shifts in their generated answers when under attack. The results in Table VI also demonstrate that our attack successfully induces PS shifts in question answering across diverse datasets. As the number of injected adversarial passages increases, the answers exhibit stronger bias, reflected by larger PS shifts across all the datasets.
Pre-trained LLMs, regardless of size or architecture, remain similarly vulnerable to bias injection attacks. Merely scaling up the model or switching architectures does not inherently improve robustness.
@@ 4.5 Ineffectiveness of Fact-Checking
To investigate whether existing automated fact-checking is a potential defense for bias injection attack, we compare both benign and adversarial passages to see if they are distinguishable with fact-checking API. Specifically, we run the WikiCheck API which return a result of “SUPPORTS” (TRUE), “REFUTES” (FALSE), or “NOT ENOUGH INFO” for each sentence in a passage.
@@
Verifiable claims in benign and adversarial passages are labeled as “SUPPORTS” or “REFUTES” by the fact-checking API [19]. Fig. 5(b) shows the distribution of support ratio (i.e., SUPPORTS / (SUPPORTS + REFUTES)). We see that adversarial passages exhibit a higher proportion of “SUPPORTS” labels. This phenomenon shows that our adversarial passages generated by LLMs use evidence that fact-checkers deem as facts.
Our attack creates passages that are as “factual” or more than benign passages to the WikiCheck fact-checker.
As a noteworthy caveat, all passages, including benign ones, contain a non-negligible fraction of “REFUTES” claims. This can arise from at least two factors. First, certain claims are only valid within the local context of the passage; when isolated and checked independently, the per-sentence fact-checker may misclassify them as false statements. Second, the current fact-checking API is not perfectly reliable and has occasional false negatives.
@@
TABLE VI: Average PS shift across different models and datasets. The proposed bias injection attack successfully biases the answers of prevalent LLM models with various parameter sizes across different datasets.
Figure 5: Fact-checking results: (a) Adversarial passages contain comparable proportions of sentences labeled as “NOT ENOUGH INFO” (0.69 and 0.68 on average for positive and negative bias injections, respectively), which are similar to that of benign passages (0.71). (b) Among verifiable claims, adversarial passages contain higher proportions of sentences labeled as “SUPPORTS” (0.73 and 0.79 on average for positive and negative bias injections, respectively) than benign passages (0.63).
5 BiasDef: A More Principled Defense
@@
We have seen that while existing defenses attenuate the attack effects, they remain suboptimal. We now present BiasDef, a post-processing method that augments existing dense retrieval pipelines to filter out adversarial passages from bias injection attackers. Rather than looking for syntactic characteristics of the attack or auxiliary fingerprints, BiasDef leverages the fact that attackers aim to craft adversarial passages that achieve high retrieval relevance, inject subtle perspective bias, and remain factual (the properties in Sec. 2.3). Sec. 6 evaluates the empirical efficacy of BiasDef.
5.1 Design Principles
An effective defense mechanism should address the attacks without making undue assumptions. Specifically:
  • Detection without priori knowledge: Although the attacker’s aim described in the properties in Section 2.3 are known, the real challenge is the lack of prior knowledge about the benign passages. If the retriever knows the exact characteristics (SS and PS) of benign passages in advance, it can easily filter out those adversarial ones by comparing their SS and PS values. But, individual benign passage may exhibit inherent bias or ideological leaning as well. One cannot always distinguish adversarial passages—intentionally crafted to skew the output—from naturally biased but benign passages, especially when both appear semantically relevant and factually correct.
  • Robustness to query and injection variability: Since the attacker can freely control both the target query and the number of injected adversarial passages, a robust defense method must consistently maintain high performance across a wide range of attack scenarios—including variations in both query content and injection intensity.
To address these challenges, BiasDef identifies and filters out adversarial passages by exploiting their skewed perspective distribution within a certain similarity range. We construct a two-dimensional feature space using the SS to the query and the PS as the two axes. Each candidate passage is plotted in this space based on its SS and PS values. Specifically, PS values are computed by performing PCA on the embeddings of the candidate passages and projecting each embedding onto the resulting polarization axis. Although the absolute values and signs of PS may differ from those computed by the attacker—since the defender does not have access to the synthetic passages used to construct the original polarization axis (see Sec. 3)—the relative PS skew described in Property 2 still holds. BiasDef identifies and filters potentially harmful content based on detected outliers in the overall passage distribution.
5.2 High-level Overview
BiasDef is a sanitization defense that, from first principles, aims to defeat the formulated attack definition given in Sec. 2.2. It uses KL-divergence which, as formalized next, is designed to exactly capture the separation of PS scores.
@@ 5.2 High-level Overview
Definition: Given a set $D_{\text{benign}}$ of benign passages and a set $D_{\text{adv}}$ of adversarial passages, we define $\mathcal{S}=\mathcal{S}(t_{ss})$ as a subset of the entire set $D_{\text{benign}}\cup D_{\text{adv}}$ that includes all passages whose SS exceeds a threshold $t_{ss}$. The gray area in Fig. 6(a) visualizes an example of such a subset. We denote the complement of this subset as $\overline{\mathcal{S}}=(D_{\text{benign}}\cup D_{\text{adv}})\setminus\mathcal{S}(t_{ss})$. Let $P_{\mathcal{S}}$ and $Q_{\mathcal{S}}$ denote the PS distributions of $\mathcal{S}$ and $\overline{\mathcal{S}}$, respectively. To facilitate computation, we discretize the PS distributions of both $\mathcal{S}$ and $\overline{\mathcal{S}}$. Let the PS axis be partitioned into $m$ equal-width bins [^8] We assume $m$ is sufficiently large (i.e., the bins are narrow enough) such that no bin contains both adversarial and benign passages when their PS values fall into disjoint intervals, as described in Property 2.. Define $c_{\mathcal{S}}(i)$ as the number of passages in $\mathcal{S}$ that fall into bin $i$, and similarly $c_{\overline{\mathcal{S}}}(i)$ for $\overline{\mathcal{S}}$. Then, the discrete PS distributions [^9] We add a small constant $\epsilon$ in piratical cases to avoid $\log(0)$ and $\log(\infty)$ when computing the KL divergence. See Appendix D for details. are computed as:
@@
Based on these definitions, we propose the following theorem that defines the security property our defense achieves.
@@
A detailed proof is provided in Appendix C. ∎
@@
Theorem-guided defense as a starting point: Theorem 1 says that any threshold $t^{*}_{ss}$ that cleanly separates $D_{\text{adv}}$ from $D_{\text{benign}}$ is guaranteed to be a local maximizer of the KL divergence $f_{\text{KL}}(t_{ss})$. Guided by this insight, our defense searches over the entire SS threshold space $t_{ss}\in[0,1]$ monotonically to identify the point where $f_{\text{KL}}(t_{ss})$ reaches a maximum away from the boundaries (i.e., excluding 0, 11):
@@
Statistical Limits: Theorem 1 gives ideal conditions for the defense to work perfectly: Attack passages, as defined in Section 3, are cleanly separable from benign. But, real-world distributions can have benign outlier passages that are statistically indistinguishable from adversarial ones (e.g., partial overlap in SS and PS distributions)—an artifact of real-world data. This poses a statistical limit for all sanitization defenses. As seen in Table I, the ideal conditions assumed by Theorem 1 are not fully met in practice, though around 75% of generated passages adhere to them (Table I). This causes false negatives and false positives. Our final procedure thus, modifies the theorem-guided defense to handle false positives and negatives in real-world data.
@@
Mitigation of false negatives: To recover false negatives—those adversarial passages with low SS values—we apply a Mahalanobis-distance-based classification [49], a classical technique widely used for identifying outlier in high-dimensional embedding spaces [50]. This method captures the covariance structure of the detected adversarial passages in $\mathcal{S}(t^{*}_{ss})$ and identifies samples that exhibit small deviations from the cluster center. Specifically, a passage $d$ is considered a false negative and is included in the refined subset if it satisfies the following condition:
$\sqrt{(E_{d}(d)-\mu)V^{-1}(E_{d}(d)-\mu)^{\top}}<T$ (11)
where $E_{d}(d)$ is the embedding of passage $d$, and $\mu$ and $V$ denote the mean vector and covariance matrix of the embeddings of passages in $\mathcal{S}(t^{*}_{ss})$, respectively. This condition evaluates the Mahalanobis distance between $d$ and the center of the adversarial cluster. The threshold $T$ is selected via hyperparameter tuning based on validation performance.
@@
After excluding the suspicious adversarial passages—those identified by first removing the false positives in $\alpha$ from $\mathcal{S}(t^{*}_{ss})$ and then adding the false negatives that satisfy Eq. (11)—from the top-$4\times k$ candidate set, we select and return the final top-$k$ relevant passages. The complete defense pseudo-code is summarized in Appendix D.
6 Evaluation of the Effectiveness of BiasDef
We now evaluate the efficacy of our aforementioned defense BiasDef against our proposed bias injection attack.
6.1 Retrieval Performance of BiasDef
Figure 7: Comparison of BiasDef with baselines in A-Recall@5 and Recall@5. Solid and hollow markers represent negative and positive bias injections, respectively. BiasDef consistently achieves the lowest A-Recall@5 while maintaining the highest Recall@5 at injection intensities of 10 and 5, and retains at least 85% of the best-performing baselines’ Recall@5 at intensities of 1 and 0.
Figure 8: CDF of BiasDef’s A-Recall@5. BiasDef completely filters out adversarial passages for over 73% queries.
@@ 6.1 Retrieval Performance of BiasDef
TABLE VII: Evading rate of adversarial passages under BiasDef’s defense. Adversarial passages that lack either Property 1 or 2 are more likely to evade the detection.
In Fig. 7, we plot BiasDef and the baselines by their average A-Recall@5 and Recall@5 across all 452 WIKI-BALANCE queries. BiasDef consistently achieves the lowest A-Recall@5 against all tested injection intensities, while maintaining nearly the highest Recall@5.
A-Recall@5: BiasDef achieves the lowest average A-Recall@5—0.16, 0.15, 0.04, 0.04, 0.15, 0.14 for #Neg=10,5,1 and #Pos=1,5,10, respectively—among all evaluated retrievers, outperforming the best baseline defense (BRRA in Table II) by an average of 15% across different injection intensities. As shown in Fig. 8, BiasDef is able to completely filter out adversarial passages from the top-5 results for approximately 73% of the queries.
We further analyze the evasion rate of adversarial passages—the proportion of injected adversarial passages that successfully bypass BiasDef’s filtering mechanism. As shown in Table VII, BiasDef is less effective at filtering out those adversarial passages that do not simultaneously satisfy both Property 1 and Property 2. This is theoretically expected too, as such atypical passages do not produce significant KL divergence changes during the defense process.
Recall@5: BiasDef achieves average Recall@5 values of 0.40, 0.41, 0.50, $5±\pm0$.04, 0.50, 0.42, and 0.41 for #Neg=10, 5, 1, the unattacked case, and #Pos=1, 5, and 10, respectively. With 5 and 10 injections, BiasDef achieves the highest Recall@5. For 0 and 1 injections, BiasDef maintains 85% of the Recall@5 achieved by No Def., the best-performing baseline. Across injection intensities, BiasDef loses only 14–38% of relevant benign passages compared to the optimum (i.e., 0.64 for No Def. in the unattacked case), achieving an average Recall@5 improvement of 62% over the best-performing baseline (SMART in Table III).
@@ 6.2 Generation Performance of BiasDef
Unattacked average absolute PS (Avg. ||PS||): BiasDef does not significantly reduce the unattacked Avg. ||PS|| compared with No Def. (0.060), indicating that BiasDef does not cause significantly new additional false positives among benign passages with strong perspectives and high PS values.
@@
6.3 Limitations, Discussion, and Future Work
Confusion Attacks on LLMs: While this paper focuses on attacks that inject adversarial passages with a consistent ideological bias, an alternative and potentially more subtle attack vector remains unexplored: deliberately retrieving a mixture of passages that present conflicting viewpoints. Such contradictory context increases PS variation in the generated answers, leading to uncertainty or ambiguity. A sophisticated adversary might exploit this by engineering “confusion attacks”—poisoning the corpus with passages that introduce semantic conflict—thereby causing the model to generate vague, indecisive, or even contradictory outputs.
Multi-dimensional extensions: Real debates can span multiple, sometimes orthogonal axes, beyond that considered in our PS score. We used the most polarizing dimension, captured by the first principal component during PCA, which typically represents the strongest binary disagreement (e.g., support vs. oppose). Nonetheless, the approach can be extended to multiple dimensions by considering additional principal components. In principle, BiasDef could be also be applied iteratively along these dimensions [51] to mitigate multi-axis polarization, but exact details remain future work.
Influence on human users: Our formulation quantifies success in retrieval and bias, but does not fully capture aspects such as answer usefulness and the attack’s influence on user trust and beliefs, which are noticeable and influence human users. A comprehensive assessment of these downstream human effects remains a challenging but vital future work.
General-purpose defense: RAG systems operate in complex environments that may be subject to multiple types of attacks, of which bias injection is only one. A promising future direction is to integrate complementary defenses (e.g., BiasDef and KAD) into a unified, general-purpose defense system capable of mitigating a broader range of threats.

### v2

@@ 5.4 Impact on the LLM Answers
Unattacked average absolute PS (Avg. ||PS||): This value, as reported in Tables 4 and 5, measures the original bias level in unattacked cases. It also indicates whether the defender misses benign passages with strong perspectives, resulting in more neutral but ambiguous content. BRRA, MMR, and SMART yield lower Avg. ||PS|| than No Def. for both retrieved passages and answers. This is because the diversity mechanism may aggressively attenuate strong perspectives.
PS shift caused by the attack: The attacker intends to make the PS of the retrieved passages biased toward its desired direction (i.e., ascending direction for positive EBI and descending direction for negative EBI). We report the average PS shift (see Eq. (4)) of the top-5 retrieved passages. The EBI attack causes the retrievers to retrieve biased passages, and their average PS shift increases as more adversarial passages are injected (see Table 4). This demonstrates the vulnerability of retrieval systems to deliberate EBI attacks through adversarial passages.
We also report on the average PS shift of the generated answers to every question in the corpus. As listed in Table 5, the EBI attack successfully induces PS shifts in the answers, and the magnitude of the shift increases with the number of injected adversarial passages. Moreover, the generated answers are positively correlated with the retrieved passages in terms of PS values. Fig. 4 quantifies such correlations.
Retrieval bias propagates through the generation process. LLM models fail to mitigate the injected biases during answer synthesis, even when equipped with critical thinking instructions in the prompt (see Prompt-Generation in Appendix A for the prompt used).
Vulnerability across models and datasets: A natural question arises: Does the impact of EBI remain if we switch to a different—or larger—LLM models? To answer this, we evaluate multiple LLMs using identical retrieval contexts, thereby isolating the model’s role in absorbing contextual bias. Table 6 presents representative results for the undefended retriever on WIKI-BALANCE.
Our results show that all three LLMs exhibit comparable PS shifts in their generated answers when under attack. The results in Table 6 also demonstrate that our attack successfully induces PS shifts in question answering across diverse datasets. As the number of injected adversarial passages increases, the answers exhibit stronger bias, reflected by larger PS shifts across all the datasets.
Pre-trained LLMs, regardless of size or architecture, remain similarly vulnerable to EBI attacks. Merely scaling up the model or switching architectures does not inherently improve robustness.
@@ 5.5 Ineffectiveness of Fact-Checking
To investigate whether existing automated fact-checking is a potential defense for EBI attacks, we compare both benign and adversarial passages to see if they are distinguishable with a standard fact-checking API. Specifically, we run the WikiCheck [15] API which return a result of “SUPPORTS” (TRUE), “REFUTES” (FALSE), or “NOT ENOUGH INFO” for each sentence in a passage.
@@
Verifiable claims in benign and adversarial passages are labeled as “SUPPORTS” or “REFUTES” by the API. Fig. 5(b) shows the distribution of support ratio (i.e., SUPPORTS / (SUPPORTS + REFUTES)). We see that adversarial passages exhibit a higher proportion of “SUPPORTS” labels. This phenomenon shows that our adversarial passages generated by LLMs use evidence that fact-checkers deem as facts.
Our attack creates passages that are as “factual" or more than benign passages to the WikiCheck fact-checker.
As a noteworthy caveat, all passages, including benign ones, contain a non-negligible fraction of “REFUTES” claims. This can arise from at least two factors. First, certain claims are only valid within the local context of the passage; when isolated and checked independently, the per-sentence fact-checker may misclassify them as false statements. Second, this can be because the current fact-checking API is not perfectly reliable and has occasional false negatives.
@@
Table 6: Average PS shift across different models and datasets. The proposed EBI attack successfully biases the answers of prevalent LLM models with various parameter sizes across different datasets.
Figure 5: Fact-checking results: (a) Adversarial passages contain comparable proportions of sentences labeled as “NOT ENOUGH INFO” (0.69 and 0.68 on average for positive and negative EBI attacks, respectively), which are similar to that of benign passages (0.71). (b) Among verifiable claims, adversarial passages contain higher proportions of sentences labeled as “SUPPORTS” (0.73 and 0.79 on average for positive and negative EBI attacks, respectively) than benign passages (0.63).
6 BiasDef: A Principled, Adaptive Defense
@@
We have seen that existing sanitizers can attenuate the impact of EBI, but remain fundamentally insufficient. Is there an adaptive defense that can defeat EBI attacks as formulated?
To explore this question, we design of a more effective defense that explicitly leverages PS-based viewpoint quantification. The rationale behind our design is as follows: Although attackers leave no explicit surface-level fingerprints, adversarial passages exhibit statistically different distributions from benign passages when observed jointly in the PS-SS space.
To test this hypothesis, we design a defense prototype, BiasDef, which relies on explainable statistical methods to filter out adversarial passages. In this process, we do not assume that adversarial and benign passages are perfectly separable; instead, we explicitly consider the more realistic case where their distributions partially overlap. The reasons are as follows. First, as shown in Table 1, while many adversarial passages exhibit extreme PS values and higher SS values than benign passages, a non-negligible fraction does not. This violates the assumption that statistical distributions of adversarial and benign passages are perfect separable. Second, note that the defender may not know the exact anchor passages used by the attacker to construct the attack. So, our proposed BiasDef derives its polarization axis from a different set of anchor passages retrieved from the corpus, which naturally induces a PS distribution that differs from that used by the attacker.
Therefore, BiasDef is composed of two components: (1) a base detector that admits a theoretical performance guarantee in the idealized case where adversarial passages are perfectly separable from benign ones (i.e., all adversarial passages satisfy Properties 1 and 2); (2) a set of statistical countermeasures that mitigate false positives and false negatives of the base detector in realistic settings where adversarial and benign distributions overlap, i.e., they are not perfectly separable.
6.1 Base Detector
BiasDef identifies and filters adversarial passages by considering the PS score within a constrained similarity range. Given a query, we first use the default retriever to obtain an expanded candidate set that contains more passages than the final top-$k$ context budget. For these candidate passages, we jointly characterize each passage by its SS to the query and its PS value, forming a two-dimensional PS-SS feature space. Specifically, BiasDef computes a polarization axis by performing PCA on the embeddings of the retrieved candidate passages, and projects each passage onto this axis to obtain its PS value. This joint representation enables BiasDef to reason about epistemic bias conditioned on relevance, rather than treating relevance and bias as independent signals. To facilitate detection, we introduce the following definition.
@@ 6.1 Base Detector
Definition: Given a set $D_{\text{benign}}$ of benign passages and a set $D_{\text{adv}}$ of adversarial passages, we define $\mathcal{S}=\mathcal{S}(t_{ss})$ as a subset of the entire set $D_{\text{benign}}\cup D_{\text{adv}}$ that includes all passages whose SS exceeds a threshold $t_{ss}$. The gray area in Fig. 6(a) visualizes an example of such a subset. We denote the complement of this subset as $\overline{\mathcal{S}}=(D_{\text{benign}}\cup D_{\text{adv}})\setminus\mathcal{S}(t_{ss})$. Let $P_{\mathcal{S}}$ and $Q_{\mathcal{S}}$ denote the PS distributions of $\mathcal{S}$ and $\overline{\mathcal{S}}$, respectively. To facilitate computation, we discretize the PS distributions of both $\mathcal{S}$ and $\overline{\mathcal{S}}$. Let the PS axis be partitioned into $m$ equal-width bins [^8] We assume $m$ is sufficiently large (i.e., the bins are narrow enough) such that no bin contains both adversarial and benign passages when their PS values fall into disjoint intervals, as described in Property 2.. Define $c_{\mathcal{S}}(i)$ as the number of passages in $\mathcal{S}$ that fall into bin $i$, and similarly $c_{\overline{\mathcal{S}}}(i)$ for $\overline{\mathcal{S}}$. Then, the discrete PS distributions [^9] We add a small constant $\epsilon$ in piratical cases to avoid $\log(0)$ and $\log(\infty)$ when computing the KL divergence. See Algorithm 1 for details. are computed as:
@@
Based on these definitions, the following theorem that defines the security property our base BiasDefdetector achieves.
@@
A detailed proof is provided in Appendix B. ∎
@@
Theorem-guided base detector: Theorem 1 says that any threshold $t^{*}_{ss}$ that cleanly separates $D_{\text{adv}}$ from $D_{\text{benign}}$ is guaranteed to be a local maximizer of the KL divergence $f_{\text{KL}}(t_{ss})$. Guided by this insight, our defense searches over the entire SS threshold space $t_{ss}\in[0,1]$ monotonically to identify the point where $f_{\text{KL}}(t_{ss})$ reaches a maximum away from the boundaries (i.e., excluding 0, 11):
@@
6.2 Mitigating False Positives and Negatives
As explained earlier, adversarial and benign passages may not be perfectly separable, i.e, some passage may be deemed both adversarial and benign with non-negligible probability. In such cases, the base detector—which is optimal in the perfect separable case—will have false positives or negatives. The following mitigation augments the base detector to deal with the possibility of false positives and negatives.
@@ 6.2 Mitigating False Positives and Negatives
Mitigation of false negatives: To recover false negatives—those adversarial passages with low SS values—we apply a Mahalanobis-distance-based classification [44], a classical technique widely used for identifying outlier in high-dimensional embedding spaces [45]. This method captures the covariance structure of the detected adversarial passages in $\mathcal{S}(t^{*}_{ss})$ and identifies samples that exhibit small deviations from the cluster center. Specifically, a passage $d$ is considered a false negative and is included in the refined subset if its embedding $E(d)$ satisfies the following condition:
$\sqrt{(E(d)-\mu)V^{-1}(E(d)-\mu)^{\top}}<T$ (12)
where $\mu$ and $V$ denote the mean vector and covariance matrix of the embeddings of passages in $\mathcal{S}(t^{*}_{ss})$, respectively. This condition evaluates the Mahalanobis distance between $d$ and the adversarial cluster. The threshold $T$ is selected via hyperparameter tuning based on validation performance.
@@
After excluding the suspicious adversarial passages—those identified by first removing the false positives in $\alpha$ from $\mathcal{S}(t^{*}_{ss})$ and then adding the false negatives that satisfy Eq. (12)—from the retrieved candidate set, we select and return the final top-$k$ relevant passages. The defense pseudo-code is summarized in in Appendix C.
6.3 Effectiveness Evaluation of BiasDef
We now evaluate the efficacy of BiasDef against our proposed attack on the same benchmark described in Sec. 5.2.
A-Recall@5: BiasDef achieves an average A-Recall@5 of 0, 0.04, 0.19, 0.20, 0.04, 0.19, 0.20 for the unattacked case, #Negative Injection = 1, 5, 10, and #Positive Injection = 1, 5, 10, respectively, which are comparable to the lowest values achieved by the baselines (BRRA in Table 2). As shown in Fig. 7, BiasDef is able to completely filter out adversarial passages from the top-5 results for over 70% of the queries.
We further analyze the evasion rate of adversarial passages—the proportion of injected adversarial passages that successfully bypass BiasDef’s filtering mechanism. As shown in Table 7, BiasDef is less effective at filtering out those adversarial passages that do not simultaneously satisfy both Property 1 and Property 2. This is theoretically expected too, as such atypical passages do not produce significant KL divergence changes during the defense process.
Figure 7: CDF of BiasDef’s A-Recall@5. BiasDef completely filters out adversarial passages for over 70% queries.
@@ 6.3 Effectiveness Evaluation of BiasDef
Table 7: Evading rate of adversarial passages under BiasDef’s defense. Adversarial passages that lack either Property 1 or 2 are more likely to evade the detection.
@@
[…]
Unattacked average absolute PS (Avg. ||PS||): BiasDef does not significantly reduce the unattacked Avg. ||PS|| (0.069) compared with No Def. (0.060), indicating that BiasDef does not falsely filter out more benign passages with strong perspectives and high PS values.
@@
7 Limitations, Discussion, and Future Work
LLM vs. Human Judgement: Our formulation provides a quantitative measure of epistemic bias. We empirically see that passages with substantially different PS values tend to express different viewpoints as deemed by an LLM acting as a judge. As per the subjective and qualitative assessment by us (authors), the PS value does correlate with "support" vs. "oppose" viewpoint and strongly biased passages have high PS value magnitudes. This observation is consistent across datasets and queries considered in our experiments.
An important open question is how PS correlate with human assessment, and how does that affect human trust, perception, and belief formation. In particular, it remains unclear what magnitude of PS shift is sufficient to induce a meaningful change in a user’s interpretation or confidence in the generated answer. A comprehensive evaluation of these downstream human effects requires controlled user studies and is beyond the scope of this paper, but represents an important direction for future work.
Multi-dimensional extensions: Real debates can span multiple, sometimes orthogonal axes, beyond that considered in our PS. We used the most polarizing dimension, captured by the first principal component during PCA, which typically represents the strongest binary disagreement (e.g., support vs. oppose). Nonetheless, the approach can be extended to multiple dimensions by considering additional principal components. In principle, BiasDef could also be applied iteratively along these dimensions [47] to mitigate multi-axis polarization, but exact details remain future work.
General-purpose defense: RAG systems operate in complex environments that may be subject to multiple types of attacks, of which EBI is only one. A promising future direction is to integrate complementary defenses (e.g., BiasDef and KAD) into a unified, general-purpose defense system capable of mitigating a broader range of threats.